OpenStudy (anonymous):

Finish writing the equation of the line y=-1/3x-? given the points (3, -5) and (-6, -2). Type only the part that is missing in the equation.

OpenStudy (anonymous):

As a general equation is y = mx+c where m is slope of the line and c is the y intercept since both the point the points are in the line, so u can plug in either (3,-5) or(-6,-2) to determine the missing link in the equation

OpenStudy (anonymous):

i get confused on where to plug them in

OpenStudy (anonymous):

it is given y = -(1/3)x -? plug them here now u can determine your missing link in the equation

OpenStudy (anonymous):

(3, -5) and (-6, -2) \[\huge y-y_1 =m(x-x_1) \] \[\huge \rightarrow y-(-5) =\frac{-5-(-2)}{3-(-6)}(x-3)\] \[\huge \rightarrow y+5) =\frac{-5+2}{3+6}(x-3)\] \[\huge \rightarrow y+5 =\frac{-3}{9}(x-3)\] \[\huge \rightarrow y+5 =-\frac{1}{3}(x-3)\] \[\huge \rightarrow y+5 =-\frac{1}{3}x+1\] \[\huge \rightarrow y =-\frac{1}{3}x+1-5\] \[\huge \rightarrow y =-\frac{1}{3}x-4\] Hence the missing part in the eq. is 4 @felicagail